# An algebraic criterion for the vanishing of bounded cohomology

### v4

@@ 5.7.1. Hamiltonian diffeomorphisms and symplectomorphisms
Let $\mathbb{R}^{2n}$ be endowed with its standard symplectic form. Let $G$ be a group such that $[\operatorname{\textup{Ham}_{c}}(\mathbb{R}^{2n}),\operatorname{\textup{Ham}_{c}}(\mathbb{R}^{2n})]\leq G\leq\operatorname{\textup{Homeo}_{c}}(\mathbb{R}^{2n})$. Then $G$ is $\operatorname{\mathfrak{X}^{\textup{sep}}}$-boundedly acyclic.
@@
We see $\mathbb{R}^{2n}$ as ××$\mathbb{R}\times\mathbb{R}\times\mathbb{R}^{2n-2}$, with coordinates $(x,y,\underline{z})$. Let $K\subset\mathbb{R}^{2n}$ be a compact set. Then there exists a compact set $A\subset\mathbb{R}^{2n-2}$ such that $K\subset B_{R}(0,0)\times A$, where $B_{R}(0,0)\subset\mathbb{R}^{2}$ is the closed disk of radius $R$ around the origin.
@@
Figure 7. The projection of the Hamiltonian vector field on the $2$-dimensional disc $B_{4R}(2R,0)$ of the first copy of $\mathbb{R}^{2}$.
@@ 5.7.2. Volume-preserving diffeomorphisms
Let $n\geq 2$ and let $\mathbb{R}^{n}$ be endowed with its standard volume form. Let $G$ be a group such that $[\operatorname{\textup{Diff}^{0}_{c,\operatorname{\textup{vol}}}}(\mathbb{R}^{n}),\operatorname{\textup{Diff}^{0}_{c,\operatorname{\textup{vol}}}}(\mathbb{R}^{n})]\leq G\leq\operatorname{\textup{Homeo}_{c}}(\mathbb{R}^{n})$. Then $G$ is $\operatorname{\mathfrak{X}^{\textup{sep}}}$-boundedly acyclic.
@@
Note that all the transformation groups encountered so far, though they have commuting cyclic conjugates, do not have commuting $\mathbb{Z}$-conjugates in general [CFFLM24, Section 3.2].
@@ 5.7.3. Portable manifolds
The key property of portable manifolds is the following:
@@ 5.7.5. Products
Let $M$ be a closed connected manifold endowed with a volume form, and let ×$\mathbb{R}\times M$ be endowed with the product volume form. Is $\operatorname{\textup{Diff}_{c,\operatorname{\textup{vol}}}}(\mathbb{R}\times M)$ ($\operatorname{\mathfrak{X}^{\textup{sep}}}$-)boundedly acyclic? In particular, is $\operatorname{\textup{Diff}_{c,\operatorname{\textup{vol}}}}(\mathbb{R}\times S^{1})$ ($\operatorname{\mathfrak{X}^{\textup{sep}}}$-)boundedly acyclic?
@@ 5.9. Highly transitive actions
Corollary 5.35 applies to Thompson’s group $T$, but also to the Stein–Thompson groups [Ste92] of the form $T_{2,n_{2},\ldots,n_{k}}$. The (circular) high transitivity is proved in [FFLZ24, Example 3.7] (note that the reference assumes that $n_{2}=3$, but this is not used in the proof of high transitivity). It also applies in other examples [FFLZ24, Examples 3.8 and 3.9] including the golden ratio Thompson’s group $T_{\tau}$ and its commutator subgroup [Cle00, BNR22] and the piecewise projective group $S$ [Lod19].
@@
[…]
@@ 5.10. Interval exchange transformations
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
The infinite length of $[0,\infty)$ plays an important role in our proof. This motivates the following question:
@@
[…]
@@ 5.11. Direct limit groups in algebraic $K$-theory
Ad (1) – (6) Let us denote by $\Gamma_{(i)}=\cup_{n\geq 1}\Gamma_{(i)}^{n}$ the groups appearing in the statement for $1\leq i\leq 5$. We are going to show that $\Gamma_{(i)}$ has commuting cyclic conjugates, so that we will conclude by Theorem 1.3. For $i=6$, we can just apply Lemma 4.11 since $[\textup{GL}(R),\textup{GL}(R)]=\textup{E}(R)$: This is called Whitehead Lemma, see for example [Mil71, Lemma 3.1].
@@
Under the above identifications, it is an element of $\Gamma_{(i)}$ for $1\leq i\leq 6$. Then $[H,{}^{t}H]=1$, as ${}^{t}H$ leaves the first $n$ basis vectors unchanged, and $[H,t^{2}]=[H,\mathrm{id}]=1$ by definition of $t$. This shows that all the groups $\Gamma_{(1)},\ldots,\Gamma_{(6)}$ have commuting cyclic conjugates and so they are $\operatorname{\mathfrak{X}^{\textup{sep}}}$-boundedly acyclic by Theorem 1.3.
@@
can be realized inside the group $\textup{O}_{2n,2n}(R)$ by exchanging the first $2n$ basis vectors with the last $2n$ basis vectors. We call this transformation $t$. Then the same argument as above shows that if for every finitely generated subgroup $H$, taking $n$ such that $H\leq\textup{O}_{n,n}(R)$, the element $t$ above satisfies $[H,{}^{t}H]=1$ and $[H,t^{2}]=[H,\operatorname{id}]=1$, so we conclude again by Theorem 1.3.
@@
Note that all the groups considered in Corollary 5.45, though they have commuting cyclic conjugates, do not have commuting $\mathbb{Z}$-conjugates in general [CFFLM24, Section 3.2].
@@ 5.12. Algebraically closed groups
In the paper were they introduced mitotic groups, Baumslag–Dyer–Heller proved that algebraically closed groups are mitotic [BDH80]. It follows that algebraically closed groups are acyclic [BDH80] and boundedly acyclic [Löh17]. We extend this to encompass non-trivial coefficients:
@@
$[g_{i},{}^{x}g_{j}]=e;\qquad[g_{i},x^{2}]=e\qquad:\qquad 1\leq i,j\leq n.$
This system is consistent, because it has a solution in the overgroup $\Gamma\wr\mathbb{Z}/2$, by setting $x$ to be the generator of $\mathbb{Z}/2$. Since $\Gamma$ is algebraically closed, there exists a solution $t\in\Gamma$. This solution must satisfy $[H,{}^{t}H]=1$ and $[H,t^{2}]=1$. This shows that $\Gamma$ has commuting cyclic conjugates and concludes the proof. ∎
@@ 5.13. The generic group
Let $\Sigma_{n}$ be the union of systems of equations and inequations that have been played by both players until stage $n-1$, and the system played by player one at stage $n$. Let $k_{n}\in\mathbb{N}$ be such that all numbers that occur as coefficients in $\Sigma_{n}$ are smaller than $k_{n}$ and, moreover, $k_{n}>n$. Let $f=k_{n}$. At this stage, player two plays the union $\Delta_{n}\cup\Delta_{n}^{\prime}$ where:
@@ 6. Bounded acyclicity and rigidity
We end with a short discussion on $\operatorname{\mathfrak{X}^{\textup{sep}}}$-bounded acyclicity and its role in rigidity theory. There are several contexts in which group actions on certain non-positively curved spaces produce characteristic classes in bounded cohomology. In such contexts, actions with trivial characteristic class are ‘elementary’ in a suitable sense, and therefore vanishing results translate into rigidity theorems: Every action must be elementary. In most classical settings, these classes belong to the second bounded cohomology with trivial real coefficients $H^{2}_{b}(-;\mathbb{R})$. This includes hyperbolic spaces [BF02, Man08, HM21], CAT(0) spaces [BF09], CAT(0) cube complexes [FFT19], Hermitian symmetric spaces [BI04, BIW07] and the circle [Ghy01]. In such contexts, weaker vanishing results suffice [Kot08, FFL23].
@@
However, group actions on the circle are the only setting in which this really leads to a complete understanding. In the other aforementioned cases, there are either strong restrictions on the types of actions to consider, or the notion of an ‘elementary’ action is quite permissive. In such cases, certain characteristic classes with $\ell^{p}$ or $L^{p}$ coefficients, $1<p<\infty$, can be used to prove more powerful rigidity theorems. This includes simply connected manifolds of pinched negative curvature [Sel92, Gro93], certain generalized notions of negatively curved spaces [MMS04, MS04, MS06] and finite-dimensional CAT(0) cube complexes [CFI16]. Up to possibly adding countability conditions on either the space or the group, the coefficients for the corresponding classes are dual and separable, and therefore we obtain rigidity results for $\operatorname{\mathfrak{X}^{\textup{sep}}}$-boundedly acyclic groups.
@@
So far we have only talked about rigidity of representations $\varrho\colon\Gamma\to G$, where $G$ is the isometry group of a non-positively curved space. One can more generally consider measurable cocycles $\sigma\colon\Gamma\times\Omega\to G$, for some probability $\Gamma$-space $\Omega$. Measurable cocycles have been central to rigidity properties for higher-rank lattices, going back to the works of Margulis and Zimmer [Zim84]. Motivated by this, several results using bounded cohomology to prove rigidity of representations, were generalized to encompass measurable cocycles [MS22a, MS22b, SS22, Sav21a, Sav21b, SS21]. The condition that ensures the rigidity of measurable cocycles is the vanishing of a characteristic class in $H^{2}_{b}(\Gamma;L^{\infty}(\Omega))$.
@@
The space $L^{\infty}(\Omega)$ is not separable, so we cannot directly apply $\operatorname{\mathfrak{X}^{\textup{sep}}}$-bounded acyclicity. However, as was noticed by Monod and Shalom [MS06, Corollary 3.5], for all countable $\Gamma$ the map $H^{2}_{b}(\Gamma;L^{\infty}(\Omega))\to H^{2}_{b}(\Gamma;L^{2}(\Omega))$ is injective. Since $L^{2}(\Omega)$ is dual and separable (because $\Omega$ is a probability measure space), we deduce that $H^{2}_{b}(\Gamma;L^{\infty}(\Omega))=0$ for every countable $\operatorname{\mathfrak{X}^{\textup{sep}}}$-boundedly acyclic group $\Gamma$, and therefore such groups enjoy the rigidity properties for measurable cocycles mentioned above.
@@
Regarding higher degrees, there are several contexts in which characteristic classes of bundles can be proved to be bounded [Fri17, Chapters 11-13]; in such contexts vanishing results for higher-degree bounded cohomology imply triviality results for spaces of bundles. Higher dimensional bounded cohomology groups also appear naturally in the study of the simplicial volume of a manifold [Fri17, Chapters 7-9]. Even if one is only interested in computing bounded cohomology in degree $2$, knowledge in degree $3$ is sometimes required to carry out computations for group extensions [Mon01, Section 12], or to compute other invariants closely related to bounded cohomology [KKM+22, GLMR23].

### v5

@@ 5.7.1. Hamiltonian diffeomorphisms and symplectomorphisms
Let $\mathbb{R}^{2n}$ be endowed with its standard symplectic form. Let $G$ be a group such that $\operatorname{\textup{Ham}_{c}}(\mathbb{R}^{2n}),\operatorname{\textup{Ham}_{c}}(\mathbb{R}^{2n})]\leq G\leq\operatorname{\textup{Homeo}_{c}}(\mathbb{R}^{2n})$. Then $G$ is $\operatorname{\mathfrak{X}^{\textup{sep}}}$-boundedly acyclic.
@@
We see $\mathbb{R}^{2n}$ as $\mathbb{R}\times\mathbb{R}\times\mathbb{R}^{2n-2}$, with coordinates $(x,y,\underline{z})$. Let $K\subset\mathbb{R}^{2n}$ be a compact set. Then there exists a compact set $A\subset\mathbb{R}^{2n-2}$ such that $K\subset B_{R}(0,0)\times A$, where $B_{R}(0,0)\subset\mathbb{R}^{2}$ is the closed disk of radius $R$ around the origin.
@@
Figure 7. The projection of the Hamiltonian vector field on the 22-dimensional disc $B_{4R}(2R,0)$ of the first copy of $\mathbb{R}^{2}$.
@@ 5.7.2. Volume-preserving diffeomorphisms
Let $n\geq 2$ and let $\mathbb{R}^{n}$ be endowed with its standard volume form. Let $G$ be a group such that $\operatorname{\textup{Diff}^{0}_{c,\operatorname{\textup{vol}}}}(\mathbb{R}^{n}),\operatorname{\textup{Diff}^{0}_{c,\operatorname{\textup{vol}}}}(\mathbb{R}^{n})]\leq G\leq\operatorname{\textup{Homeo}_{c}}(\mathbb{R}^{n})$. Then $G$ is $\operatorname{\mathfrak{X}^{\textup{sep}}}$-boundedly acyclic.
@@
Note that all the transformation groups encountered so far, though they have commuting cyclic conjugates, do not have commuting $\mathbb{Z}$-conjugates in general [CFFLM25, Section 3.2].
@@ 5.7.3. Portable manifolds
Any manifold that splits as the direct product of a closed manifold and $\mathbb{R}^{n}$ is portable, on the other hand compact manifolds cannot be portable. The key property of portable manifolds is the following:
@@ 5.7.5. Products
Let $M$ be a closed connected manifold endowed with a volume form, and let $\mathbb{R}\times M$ be endowed with the product volume form. Is $\operatorname{\textup{Diff}_{c,\operatorname{\textup{vol}}}}(\mathbb{R}\times M)$ ($\operatorname{\mathfrak{X}^{\textup{sep}}}$-)boundedly acyclic? In particular, is $\operatorname{\textup{Diff}_{c,\operatorname{\textup{vol}}}}(\mathbb{R}\times S^{1})$ ($\operatorname{\mathfrak{X}^{\textup{sep}}}$-)boundedly acyclic?
@@ 5.9. Highly transitive actions
Corollary 5.35 applies to Thompson’s group $T$, but also to the Stein–Thompson groups [Ste92] of the form $T_{2,n_{2},\ldots,n_{k}}$. The (circular) high transitivity is proved in [FFLZ24b, Example 3.7] (note that the reference assumes that $n_{2}=3$, but this is not used in the proof of high transitivity). It also applies in other examples [FFLZ24b, Examples 3.8 and 3.9] including the golden ratio Thompson’s group $T_{\tau}$ and its commutator subgroup [Cle00, BNR22] and the piecewise projective group $S$ [Lod19].
@@
$\cdots\to\mathbb{R}\xrightarrow{\textup{id}}\mathbb{R}\xrightarrow{0}\mathbb{R}\to\cdots$
@@ 5.10. Interval exchange transformations
Recall that an interval exchange transformation of the half-open interval $[0,1)$ is a right-continuous bijection given by cutting $[0,1)$ into finitely many half-open subintervals and rearranging by translations. We denote by $\operatorname{\textup{IET}}([0,1))$ the group of interval exchange transformations of $[0,1)$. We can define analogously the group $\operatorname{\textup{IET}}([0,\infty))$, where again we cut into finitely many half-open subintervals.
@@
$\operatorname{\textup{IET}}([0,\infty))$ is the direct union of the groups $\{\operatorname{\textup{IET}}([0,n))\}_{n\geq 1}$. In particular, $\operatorname{\textup{IET}}([0,\infty))$ is amenable, respectively contains non-abelian free subgroups, if and only if $\operatorname{\textup{IET}}([0,1))$ does.
@@
The first statement follows from the fact that the unique half-open interval of infinite length has to be preserved by an $\operatorname{\textup{IET}}$. The second statement follows from the isomorphism $\operatorname{\textup{IET}}([0,1))\cong\operatorname{\textup{IET}}([0,n))$ given by affine conjugacy, and the fact that direct unions preserve amenability and the property of not containing free subgroups. ∎
To our knowledge, nothing is known about the bounded cohomology of $\operatorname{\textup{IET}}([0,1))$. For the half line, we obtain:
@@
The group $\operatorname{\textup{IET}}([0,\infty))$ is $\operatorname{\mathfrak{X}^{\textup{sep}}}$-boundedly acyclic, and so are its derived subgroups.
@@
By Lemma 5.41, a finitely generated subgroup $H\leq\operatorname{\textup{IET}}([0,\infty))$ is contained in $\operatorname{\textup{IET}}([0,n))$ for some $n\geq 1$. Let $t$ be the $\operatorname{\textup{IET}}$ exchanging $[0,n)$ with $[n,2n)$. Then $^{t}H\leq\operatorname{\textup{IET}}([n,2n))$ and thus $H,{}^{t}H]=1$. Moreover $H,t^{2}]=[H,1]=1$, so $\operatorname{\textup{IET}}([0,\infty))$ has commuting cyclic conjugates, and we conclude by Theorem 1.3. The statement on derived subgroups follows from Lemma 4.11. ∎
@@
The infinite length of $0,\infty)$ plays an important role in our proof. This motivates the following question:
@@
Is the group $\operatorname{\textup{IET}}([0,1))$ ($\operatorname{\mathfrak{X}^{\textup{sep}}}$-)boundedly acyclic?
Note that a negative answer would imply that $\operatorname{\textup{IET}}([0,1))$ is non-amenable.
@@ 5.11. Direct limit groups in algebraic $K$-theory
Ad (1) – (6) Let us denote by $\Gamma_{(i)}=\cup_{n\geq 1}\Gamma_{(i)}^{n}$ the groups appearing in the statement for $1\leq i\leq 5$. We are going to show that $\Gamma_{(i)}$ has commuting cyclic conjugates, so that we will conclude by Theorem 1.3. For $i=6$, we can just apply Lemma 4.11 since $\textup{GL}(R),\textup{GL}(R)]=\textup{E}(R)$: This is called Whitehead Lemma, see for example [Mil71, Lemma 3.1].
@@
Under the above identifications, it is an element of $\Gamma_{(i)}$ for $1\leq i\leq 6$. Then $H,{}^{t}H]=1$, as $^{t}H$ leaves the first $n$ basis vectors unchanged, and $H,t^{2}]=[H,\mathrm{id}]=1$ by definition of $t$. This shows that all the groups $\Gamma_{(1)},\ldots,\Gamma_{(6)}$ have commuting cyclic conjugates and so they are $\operatorname{\mathfrak{X}^{\textup{sep}}}$-boundedly acyclic by Theorem 1.3.
@@
can be realized inside the group $\textup{O}_{2n,2n}(R)$ by exchanging the first $2n$ basis vectors with the last $2n$ basis vectors. We call this transformation $t$. Then the same argument as above shows that if for every finitely generated subgroup $H$, taking $n$ such that $H\leq\textup{O}_{n,n}(R)$, the element $t$ above satisfies $H,{}^{t}H]=1$ and $H,t^{2}]=[H,\operatorname{id}]=1$, so we conclude again by Theorem 1.3.
@@
Note that all the groups considered in Corollary 5.45, though they have commuting cyclic conjugates, do not have commuting $\mathbb{Z}$-conjugates in general [CFFLM25, Section 3.2].
@@ 5.12. Algebraically closed groups
In the paper where they introduced mitotic groups, Baumslag–Dyer–Heller proved that algebraically closed groups are mitotic [BDH80]. It follows that algebraically closed groups are acyclic [BDH80] and boundedly acyclic [Löh17]. We extend this to encompass non-trivial coefficients:
@@
$g_{i},{}^{x}g_{j}]=e;\qquad[g_{i},x^{2}]=e\qquad:\qquad 1\leq i,j\leq n.$
This system is consistent, because it has a solution in the overgroup $\Gamma\wr\mathbb{Z}/2$, by setting $x$ to be the generator of $\mathbb{Z}/2$. Since $\Gamma$ is algebraically closed, there exists a solution $t\in\Gamma$. This solution must satisfy $H,{}^{t}H]=1$ and $H,t^{2}]=1$. This shows that $\Gamma$ has commuting cyclic conjugates and concludes the proof. ∎
@@ 5.13. The generic group
Let $\Sigma_{n}$ be the union of systems of equations and inequations that have been played by both players until stage n1n-1, and the system played by player one at stage $n$. Let $k_{n}\in\mathbb{N}$ be such that all numbers that occur as coefficients in $\Sigma_{n}$ are smaller than $k_{n}$ and, moreover, $k_{n}>n$. Let $f=k_{n}$. At this stage, player two plays the union $\Delta_{n}\cup\Delta_{n}^{\prime}$ where:
@@ 6. Bounded acyclicity and rigidity
We end with a short discussion on $\operatorname{\mathfrak{X}^{\textup{sep}}}$-bounded acyclicity and its role in the study of representations $\varrho\colon\Gamma\to G$, where $G$ is the isometry group of a non-positively curved space $X$. Recall that associated with each representation $\varrho\colon\Gamma\to G$ there is an induced action of $\Gamma$ on the space $X$. Such actions produce bounded characteristic classes in the bounded cohomology of $\Gamma$ and actions with a trivial bounded characteristic class are ‘elementary’ in a suitable sense. More precisely, the vanishing of the bounded characteristic classes of $\Gamma$ yields a rigidity theorem: All the actions induced by a representation $\varrho\colon\Gamma\to G$ must be elementary. In most classical settings, these bounded cohomology classes belong to the second bounded cohomology with trivial real coefficients $H^{2}_{b}(-;\mathbb{R})$. This happens, e.g., when $X$ is one of the following non-positively curved spaces: Hyperbolic spaces [BF02, Man08, HM21], CAT(0) spaces [BF09], CAT(0) cube complexes [FFT19], Hermitian symmetric spaces [BI04, BIW07] and the circle [Ghy01, Mat86]. In such contexts, also weaker vanishing results suffice [Kot08, FFL23].
@@
However, group actions on the circle are the only setting in which this really leads to a complete understanding. In the other aforementioned cases, there are either strong restrictions on the types of actions to consider, or the notion of an ‘elementary’ action is quite permissive. In such cases, certain characteristic classes with $\ell^{p}$ or $L^{p}$ coefficients, $1<p<\infty$, can be used to prove more powerful rigidity theorems. Spaces for which stronger rigidity results can be obtained include simply connected manifolds of pinched negative curvature [Sel92, Gro93], certain generalized notions of negatively curved spaces [MMS04, MS04, MS06] and finite-dimensional CAT(0) cube complexes [CFI16]. Up to possibly adding countability conditions on either the space or the group, the coefficients for the corresponding classes are dual and separable, and therefore we obtain rigidity results for $\operatorname{\mathfrak{X}^{\textup{sep}}}$-boundedly acyclic groups.
@@
So far we have only talked about rigidity of representations. One can more generally consider measurable cocycles $\sigma\colon\Gamma\times\Omega\to G$, for some probability $\Gamma$-space $\Omega$. Measurable cocycles have been central to rigidity properties for higher-rank lattices, going back to the works of Margulis and Zimmer [Zim84]. Motivated by this, several results using bounded cohomology to prove rigidity of representations, were generalized to encompass measurable cocycles [MS22a, MS22b, SS22, Sav21a, Sav21b, SS23]. The condition that ensures the rigidity of measurable cocycles is the vanishing of a characteristic class in $H^{2}_{b}(\Gamma;L^{\infty}(\Omega))$.
@@
The space $L^{\infty}(\Omega)$ is not separable, so we cannot directly apply $\operatorname{\mathfrak{X}^{\textup{sep}}}$-bounded acyclicity. However, as was noticed by Monod and Shalom [MS06, Corollary 3.5], for all countable $\Gamma$ the map $H^{2}_{b}(\Gamma;L^{\infty}(\Omega))\to H^{2}_{b}(\Gamma;L^{2}(\Omega))$ is injective. Since $L^{2}(\Omega)$ is dual and separable (because $\Omega$ is a probability measure space), we deduce that $H^{2}_{b}(\Gamma;L^{\infty}(\Omega))=0$ for every countable $\operatorname{\mathfrak{X}^{\textup{sep}}}$-boundedly acyclic group $\Gamma$, and therefore such groups enjoy the rigidity properties for measurable cocycles mentioned above. [^1] Since the first version of this paper, our vanishing results have been extended to encompass dual semi-separable coefficients (in the sense of Monod [Mon10]), see [Bog25]). This includes in particular the spaces $L^{\infty}(\Omega)$ discussed here.
@@
Regarding higher degrees, there are several contexts in which characteristic classes of bundles can be proved to be bounded [Fri17, Chapters 11-13]; in such contexts vanishing results for higher-degree bounded cohomology imply triviality results for spaces of bundles. Higher dimensional bounded cohomology groups also appear naturally in the study of the simplicial volume of a manifold [Fri17, Chapters 7-9]. Even if one is only interested in computing bounded cohomology in degree $2$, knowledge in degree $3$ is sometimes required to carry out computations for group extensions [Mon01, Section 12], or to compute other invariants closely related to bounded cohomology [KKM+24, GLMR23].